OpenStudy (anonymous):

to prove A=B, can i multiply/add a constant on both sides in order to reduce both LHS and RHS to simpler expressions?

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

as in a stupid example, if i want to prove x+2/2=x+1 can i do (x+2/2)-1 = (x+1)-1 x = x

ganeshie8 (ganeshie8):

yes looks good to me

ganeshie8 (ganeshie8):

the equality is not changed when you add/remove same thing both sides

OpenStudy (anonymous):

but in a proof situation that they want you to prove A=B, in your proof steps are you allowed to do for example A/2=B/2, in order to simplify things

OpenStudy (anonymous):

@satellite73

OpenStudy (loser66):

I think you should post your problem to easily observe.